To what amount will the following investment accumulate?
$462, invested today for 21 years at 16.44 percent, compounded monthly.
Answer #1

Future value = present value * (1+r)^n

Present value = $462

r = interest rate per period = 16.44%/12

n = number of periods = 21 years = 21*12 = 252 Months

=>

future value = 462 * (1+16.44%/12)^252

= $14,250.25